The potential energy of a particle executing SHM alon x-axis is U=U0(1-cosax).What is the frequency of oscillation?

Grade:12

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student

U=U0(1-cosax)

Force=-dU/dx=-aU0sinax=-w^2 (x)

from here you can easliy calculate frequency on comparing the two..

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

U = U0(1-cosax)

F = -dU/dx = -Uoasinax

for small x , sinax is small or sinax = ax

F = -U0a2x

a = -Uoa2x/m                              (a = -w2x)

-w2x = -Uoa2x/m

 w = (Uoa2/m)1/2

vikas askiitian expert
13 years ago
U = U0(1-cosax) F = -dU/dx = -Uoasinax for small x , sinax is small or sinax = ax F = -U0a2x a = -Uoa2x/m (a = -w2x) -w2x = -Uoa2x/m w = (Uoa2/m)1/2